# Charting the Landscape of Nefarious Uses of Generative Artificial Intelligence for Online Election Interference

### v2

@@ Introduction
The advent of GenAI is revolutionizing fields from human-AI teaming to creative content generation, enabling new forms of human-computer interaction and augmenting human capabilities in unprecedented ways (Brown et al., 2020; OpenAI, 2023; Rahwan et al., 2019). However, these advancements come with significant risks: here we focus on the problem of interference with democracy (Bovet and Makse, 2019; Ozmen Garibay et al., 2023) within the broader context of AI ethics (Independent High-Level Expert Group on Artificial Intelligence, 2019; Mittelstadt et al., 2016).
The primary objective of this study is to investigate the nefarious applications of GenAI in online election interference, exploring how these technologies can disrupt democratic processes. We aim to provide a comprehensive overview of the threats posed by GenAI, supported by a dual approach that includes a literature review and analysis to identify key trends and case studies.
Online election interference refers to the manipulation of digital information and communication technologies to influence the outcome of elections. This can include activities such as spreading misinformation, creating fake news, and orchestrating social media campaigns to sway public opinion. The introduction of GenAI into this arena has magnified these threats by enabling more sophisticated and scalable forms of interference (Ferrara, 2024a; Mozes et al., 2023).
@@
Generative AI, with its ability to create realistic text, images, and videos, presents new opportunities for malicious actors to deceive and manipulate the public. For example, deepfake technology can generate realistic videos of political figures saying or doing things they never did, eroding public trust in authentic information sources (see Fig. 1) (Seymour et al., 2023). Moreover, the widespread use of AI-powered botnets can amplify misinformation and create the illusion of widespread support or opposition to certain political ideas, further distorting the democratic process (Yang and Menczer, 2023).
@@
In our analysis, we identified several key case studies that illustrate the impact of GenAI on election interference. For instance, during the 2016 U.S. presidential election, AI-generated botnets played a significant role in spreading misinformation and influencing public opinion (Ferrara et al., 2016). Similarly, deepfake videos have been used to create false narratives about political candidates, affecting voter perceptions and trust (Seymour et al., 2023).
The deployment of GenAI in online election interference poses a unique challenge to democracies worldwide. Traditional methods of combating misinformation and digital manipulation are often insufficient to address the sophisticated techniques enabled by GenAI. This necessitates a comprehensive understanding of the various ways in which GenAI can be exploited for nefarious purposes and the development of new strategies to mitigate these risks (Ferrara, 2024a; Ziems et al., 2023).
@@
This perspective article aims to shed light on the darker applications of GenAI in the context of online election interference. By examining the various nefarious uses of these technologies, we hope to highlight the urgent need for regulatory oversight, technological solutions, public awareness, and collaborative efforts to safeguard the integrity of democratic processes.
Through this exploration, we seek to provide a foundation for policymakers, technologists, and civil society to develop robust strategies to protect elections from AI-driven interference, ensuring that the benefits of GenAI are not overshadowed by its potential for harm (Rudin, 2019; Cao et al., 2023).
@@
GenAI and LLMs, celebrated for their ability to process natural language and generate multimodal content, offer numerous benefits, including enhancing creativity and solving complex problems (Brown et al., 2020; OpenAI, 2023). These technologies have enabled advancements in various fields, from natural language understanding and translation to creative industries such as art and music (Cao et al., 2023). However, these same capabilities can be exploited for malicious purposes, such as creating deepfakes, synthetic identities, and targeted misinformation campaigns (Mozes et al., 2023).
@@
On the positive side, GenAI can assist in drafting documents, writing code, generating realistic images and videos, and even conducting complex data analysis (Ziems et al., 2023). These applications have immense potential to boost productivity and innovation across different sectors. For example, AI-generated content can support marketing efforts by creating personalized advertisements or generating engaging social media posts. In healthcare, GenAI can help in diagnosing diseases by analyzing medical images or generating detailed patient reports (Davenport and Kalakota, 2019).
Despite these benefits, the dark side of GenAI cannot be ignored. The same technology that generates realistic images and videos for legitimate purposes can also create deepfakes, which are manipulated media that convincingly alter appearances or actions of individuals. Deepfakes can be used to spread false information, damage reputations, and even blackmail individuals (Seymour et al., 2023). The ease with which these deepfakes can be created poses a significant threat to the authenticity of information shared online.
Similarly, LLMs can be used to generate convincing fake news articles, social media posts, and even entire personas (Chen and Shu, 2023; Augenstein et al., 2024; Choi and Ferrara, 2024a, b; Quelle and Bovet, 2024). These synthetic identities can be employed to infiltrate social networks, manipulate public opinion, and disrupt democratic processes (Haider et al., 2023; Luceri et al., 2024). AI-generated text can mimic human writing so effectively that distinguishing between genuine and fake content becomes challenging (Zellers et al., 2019; Augenstein et al., 2024; Quelle and Bovet, 2024; Jones and Bergen, 2024).
Moreover, GenAI’s ability to tailor messages to specific demographics makes it a powerful tool for targeted misinformation campaigns. These campaigns can be designed to exploit societal divisions, amplify existing biases, and influence voter behavior (Vosoughi et al., 2018). The targeted nature of these campaigns increases their effectiveness and the potential harm they can cause.
@@
The dual nature of GenAI highlights the need for a balanced approach in leveraging these technologies. While the benefits are substantial, the risks are equally significant (Ferrara, 2024b). As we continue to develop and deploy GenAI, it is crucial to implement safeguards that prevent its misuse (Rudin, 2019). This includes developing robust detection mechanisms, establishing ethical guidelines, and promoting public awareness about the potential dangers of AI-generated content (Ferrara, 2024a).
In conclusion, while GenAI holds great promise for enhancing various aspects of society, it also presents substantial risks. Understanding the dual nature of these technologies is essential for developing strategies that maximize their benefits while mitigating their potential harms (Seymour et al., 2023; Mozes et al., 2023; Ferrara, 2024b).
@@
GenAI can also be used to create convincing synthetic identities and fake accounts, which can be employed to infiltrate online communities, spread misinformation, and gather intelligence on political opponents (Ferrara, 2023b, 2024a). These synthetic identities can come with complete profiles, including photos, bios, and posting histories, making them difficult to distinguish from real users (Westerlund, 2019). The creation of fake accounts can facilitate a range of malicious activities. For instance, these accounts can be used to spread false narratives, engage in coordinated attacks against political figures, or even influence online polls. The use of synthetic identities to manipulate online discourse presents a significant challenge for social media platforms and requires sophisticated detection and countermeasures (Haider et al., 2023).
@@ Conclusions
Methodology
To achieve our research objectives, we employed a dual methodology comprising a comprehensive literature review and analysis. The literature review involved systematically identifying and analyzing scholarly articles, reports, and case studies related to the use of GenAI in online election interference. We focused on sources that provide empirical evidence and theoretical insights into the nefarious applications of these technologies. The analysis was conducted to identify key trends, influential works, and research gaps in the field. We analyzed recent journal and conference publication patterns, for a comprehensive understanding of the topic’s evolution.

### v3

@@ Introduction
The advent of GenAI is revolutionizing fields from human-AI teaming to creative content generation, enabling new forms of human-computer interaction and augmenting human capabilities in unprecedented ways (Brown et al., 2020; OpenAI, 2023; Rahwan et al., 2019). However, these advancements come with significant risks: here we focus on the problem of interference with democracy (Bovet and Makse, 2019; Ozmen Garibay et al., 2023) within the broader context of AI ethics (Independent High-Level Expert Group on Artificial Intelligence, 2019; Mittelstadt et al., 2016). The primary objective of this study is to investigate the nefarious applications of GenAI in online election interference, exploring how these technologies can disrupt democratic processes. We aim to provide a comprehensive overview of the threats posed by GenAI, illustrating dual nature paradox that this technology poses, by means of a literature review and analysis to identify key trends and case studies (Barrett et al., 2023).
@@
Online election interference refers to the manipulation of digital information and communication technologies to influence the outcome of elections (Tenove et al., 2018; Kreps and Kriner, 2023). This can include activities such as spreading misinformation, creating fake news, and orchestrating social media campaigns to sway public opinion. The introduction of GenAI into this arena has magnified these threats by enabling more sophisticated and scalable forms of interference (Goldstein et al., 2023; Mozes et al., 2023; Shoaib et al., 2023). Generative AI, with its ability to create realistic text, images, and videos, presents new opportunities for malicious actors to deceive and manipulate the public. For example, deepfake technology can generate realistic videos of political figures saying or doing things they never did, eroding public trust in authentic information sources (see Fig. 1) (Seymour et al., 2023; Barrett et al., 2023; Kenthapadi et al., 2023; Lorenz et al., 2023; Gabriel et al., 2024). Moreover, the widespread use of AI-powered botnets can amplify misinformation and create the illusion of widespread support or opposition to certain political ideas, further distorting the democratic process (Unver, 2018; Yang and Menczer, 2023; Kreps and Kriner, 2023; Coeckelbergh, 2023).
@@
In our review, we identified several key case studies that illustrate the impact of GenAI on election interference (Yang et al., 2019; Kreps et al., 2022; Chennupati, 2024). For instance, during the 2016 U.S. presidential election, AI-generated botnets played a significant role in spreading misinformation and influencing public opinion (Bessi and Ferrara, 2016; Howard et al., 2018; Shao et al., 2018; Boichak et al., 2018). Similarly, deepfake videos have been used to create false narratives about political candidates, affecting voter perceptions and trust (Vaccari and Chadwick, 2020; Dobber et al., 2021; Diakopoulos and Johnson, 2021; Seymour et al., 2023). The deployment of GenAI in online election interference poses a unique challenge to democracies worldwide. Traditional methods of combating misinformation and digital manipulation are often insufficient to address the sophisticated techniques enabled by GenAI. This necessitates a comprehensive understanding of the various ways in which GenAI can be exploited for nefarious purposes and the development of new strategies to mitigate these risks (Xu et al., 2023; Gupta et al., 2023; Ziems et al., 2023).
@@
This perspective article aims to shed light on the darker applications of GenAI in the context of online election interference. By examining the various nefarious uses of these technologies, we hope to highlight the urgent need for regulatory oversight, technological solutions, public awareness, and collaborative efforts to safeguard the integrity of democratic processes. Through this exploration, we seek to provide a foundation for policymakers, technologists, and civil society to develop robust strategies to protect elections from AI-driven interference, ensuring that the benefits of GenAI are not overshadowed by its potential for harm (Rudin, 2019; Cao et al., 2023).
@@
GenAI and LLMs, celebrated for their ability to process natural language and generate multimodal content, offer numerous benefits, including enhancing creativity and solving complex problems (Brown et al., 2020; OpenAI, 2023; Shen et al., 2024; Yao et al., 2024). These technologies have enabled advancements in various fields, from natural language understanding and translation to creative industries such as art and music (Cao et al., 2023; Franceschelli and Musolesi, 2023). However, these same capabilities can be exploited for malicious purposes, such as creating deepfakes, synthetic identities, and targeted misinformation campaigns (Barrett et al., 2023; Mozes et al., 2023; Kang et al., 2024).
@@
On the positive side, GenAI can assist in drafting documents, writing code, generating realistic images and videos, and even conducting complex data analysis (Mandapuram et al., 2018; Bail, 2024; Li et al., 2024). These applications have immense potential to boost productivity and innovation across different sectors (Ooi et al., 2023; Kanbach et al., 2024). For example, AI-generated content can support marketing efforts by creating personalized advertisements or generating engaging social media posts. In healthcare, GenAI can help in diagnosing diseases by analyzing medical images or generating detailed patient reports (Davenport and Kalakota, 2019; Sai et al., 2024).
Despite these benefits, the dark side of GenAI cannot be ignored (Wach et al., 2023; Kang et al., 2024). The same technology that generates realistic images and videos for legitimate purposes can also create deepfakes, which are manipulated media that convincingly alter appearances or actions of individuals. Deepfakes can be used to spread false information, damage reputations, and even blackmail individuals (Vaccari and Chadwick, 2020; Diakopoulos and Johnson, 2021; Seymour et al., 2023). The ease with which these deepfakes can be created poses a significant threat to the authenticity of information shared online. Similarly, LLMs can be used to generate convincing fake news articles, social media posts, and even entire personas (Chen and Shu, 2023; Augenstein et al., 2024; Choi and Ferrara, 2024a, b; Quelle and Bovet, 2024). These synthetic identities can be employed to infiltrate social networks, manipulate public opinion, and disrupt democratic processes (Haider et al., 2023; Luceri et al., 2024). AI-generated text can mimic human writing so effectively that distinguishing between genuine and fake content becomes challenging (Zellers et al., 2019; Augenstein et al., 2024; Quelle and Bovet, 2024; Jones and Bergen, 2024). Moreover, GenAI’s ability to tailor messages to specific demographics makes it a powerful tool for targeted misinformation campaigns. These campaigns can be designed to exploit societal divisions, amplify existing biases, and influence voter behavior (Boichak et al., 2018; Sap et al., 2019; Bovet and Makse, 2019; Goldstein et al., 2023; Ezzeddine et al., 2023; Hofmann et al., 2024). The targeted nature of these campaigns increases their effectiveness and the potential harm they can cause.
@@
The dual nature of GenAI highlights the need for a balanced approach in leveraging these technologies. While the benefits are substantial, the risks are equally significant (Weidinger et al., 2021, 2022; Barrett et al., 2023; Ferrara, 2024b). As we continue to develop and deploy GenAI, it is crucial to implement safeguards that prevent its misuse (Rudin, 2019). This includes developing robust detection mechanisms, establishing ethical guidelines, and promoting public awareness about the potential dangers of AI-generated content (Mittelstadt et al., 2016; Shahriari and Shahriari, 2017; Independent High-Level Expert Group on Artificial Intelligence, 2019; Jobin et al., 2019; Smuha, 2019; Yang et al., 2019; Qiu et al., 2023; Zhou et al., 2023; Sison et al., 2023). In conclusion, while GenAI holds great promise for enhancing various aspects of society, it also presents substantial risks. Understanding the dual nature of these technologies is essential for developing strategies that maximize their benefits while mitigating their potential harms (Seymour et al., 2023; Mozes et al., 2023).
Survey Approach
To sift through the growing scholarly work about the risks of GenAI, we carried out a systematic literature review and analysis. The literature review involved systematically identifying and analyzing scholarly articles, reports, and case studies related to the use of GenAI in online election interference. We focused on sources that provide empirical evidence and theoretical insights into the nefarious applications of these technologies. The analysis was conducted to identify key trends, influential works, and research gaps in the field. We analyzed recent journal and conference publication patterns, for a comprehensive understanding of the topic’s evolution.
@@
GenAI can also be used to create convincing synthetic identities and fake accounts, which can be employed to infiltrate online communities, spread misinformation, and gather intelligence on political opponents (Ferrara, 2023b, 2024a). These synthetic identities can come with complete profiles, including photos, bios, and posting histories, making them difficult to distinguish from real users (Westerlund, 2019; Ezzeddine et al., 2023). The creation of fake accounts can facilitate a range of malicious activities. For instance, these accounts can be used to spread false narratives, engage in coordinated attacks against political figures, or even influence online polls. The use of synthetic identities to manipulate online discourse presents a significant challenge for social media platforms and requires sophisticated detection and countermeasures (Haider et al., 2023).